# Product Design Using a Generative Adversarial Network: Incorporating Consumer Preferences and External Data

### v2

@@ 5.3 Performance of the Generative Models
Note: This table presents the average choice probabilities predicted for real images and images generated using different GAN models. The first (second) row shows the choice probabilities predicted by Prediction Model 1 (Prediction Model 2). We normalize the first column to 1 (i.e., make the choice probability of the real image equal 1) and present the other columns as percentage improvements relative to the first column.
In addition to the average predicted choice probabilities in Table 3, we also present the histograms of the predicted choice probabilities for the internal real images, DCGAN-generated images, and CcGAN-generated images in Figure 12. Specifically, Figure 12 (a) shows the choice probabilities predicted by Prediction Model 1, and Figure 12 (b) shows the choice probabilities predicted by Prediction Model 2. The findings from the histograms further confirm the findings in Table 3. First, incorporating consumer preference information is helpful because the CcGAN always outperforms the DCGAN. Second, incorporating external consumer preferences is particularly helpful. The CcGAN-generated images are not significantly better than the real internal images if only internal preferences are incorporated. However, when both internal and external preferences are added, the CcGAN-generated images are significantly better than the real internal images.
Overall, the results highlight the importance of incorporating internal and external preference information. The external preference information is particularly powerful. It allows the predictor model to fully recognize the external features that are potentially appealing to the internal consumers, thereby guiding the generation process to fully incorporate these attractive external features.
Figure 12: Histograms of the Predicted Choice Probabilities
Refer to caption Refer to caption
(a) Predicted by Prediction Model 1 (b) Predicted by Prediction Model 2
Note: This figure plots the histograms of the predicted average choice probabilities for real images and images generated using different GAN models.
Performance based on outside metrics
The previous performance test used a prediction model within our framework to show the popularity of the generated images. We further use metrics generated outside our framework to evaluate the popularity of the generated images. Metrics outside our framework can serve as more objective and independent measures, providing further support for our results.
We propose a distance-based metric as an indicator of popularity. Intuitively, popular internal templates have common features, suggesting that consumers may favor certain types of templates. Therefore, we can look at how similar (far way) the generated images are to the popular (unpopular) internal templates. The more closely the newly generated templates resemble the known popular templates, the more likely that they are also appealing to consumers.
The calculation procedure of the distance-based metric is as follows. First, we identify the three most-preferred internal templates $\mathcal{J}=\{j_{1},j_{2},j_{3}\}$ and the three least-preferred templates $\mathcal{K}=\{k_{1},k_{2},k_{3}\}$ from the observed consumer choices. Second, we calculate a 1000-dimensional vector embedding for the six identified internal templates and the new templates generated by different generative models. Third, we calculate the distance between each generated template $i$ and the most- and least-preferred internal templates. We further normalize the distances to be within the range of [0, 1]. Let $d_{ij}$ and $d_{ik}$ denote the normalized distances between template $i$ and the most-preferred template $j$ and the least-preferred template $k$, respectively. Finally, we calculate the distance-based metric for each generated template and average the results over all templates generated from the same generative model to obtain a measure for each generative model. To summarize, the distance-based metric for each generative model is defined as
where $N$ denotes the total number of templates generated by each generative model. This distance-based metric decreases when the generated templates are more (less) similar to the most-preferred (least-preferred) internal templates. The smaller the distance-based metric is, the better. The metric provides an independent assessment of model performance because it does not rely on the prediction model within our framework. Instead, it leverages the intrinsic similarity among the image features to suggest popularity.
Table 4 presents the distance-based metric results produced by different generative models. We find that the distance metric is smaller for the CcGAN with internal preferences and smallest for the CcGAN with both internal and external preferences. The results suggest the value of incorporating internal and external preferences into the process of generating new templates. Figure 13 provides histograms for the distance-based metrics yielded by different generative models. This result further confirms that the CcGAN with both internal and external preferences performs best.
Table 4: Distance-Based Metrics Produced for the Generated Images
@@
Statistics DCGAN CcGAN (Internal) CcGAN (Internal and External)
Mean 0.58 0.54 0.44
Median 0.55 0.52 0.40
@@
Note: This table presents the mean and median values of the distance-based metrics produced for images generated using different methods. A small value means that the generated templates are close to the popular internal templates while far away from the unpopular internal templates in terms of visual characteristics. A small metric suggests a higher degree of popularity among consumers.
Figure 13: Histograms of the Distance-Based Metrics
Note: This figure plots the histograms of the distance-based metrics for the images generated using different GAN models. A small value means that the generated templates are close to the popular internal templates while far away from the unpopular internal templates in terms of visual characteristics. A small metric suggests a higher degree of popularity among consumers.
6 Managerial Implications and Conclusion
In this paper, we provide a practical solution for the product design decisions of firms using state-of-the-art deep learning methods (i.e., GANs). The unique contribution of our work is that our proposed model systematically incorporates consumer preference information and external UGC information into the design generation process. First, we construct popularity labels based on a prediction model trained on historical consumer choices and include the labels as additional inputs to guide the training process of the generative model. Such an approach enables the generative model to directionally generate new product designs that are more likely to be preferred by consumers. Second, our model allows firms to leverage the rich design ideas and consumer tastes contained in UGC. Both the predictor model and the generative model are trained on external product designs and external consumer preferences. The external images not only serve as extra training resources for the generative models but also provide additional external consumer preference information that can be utilized to guide the generative model.
Two important managerial takeaways can be derived from our results. First, it is beneficial to incorporate consumer preference information into the generation process because it can significantly improve the popularity and quality of the generated designs. Second, it is not sufficient to only incorporate external designs; external preference information is also required for the generative model to fully incorporate new appealing features from external designs. When the predictor model is trained only on internal preferences, consumers cannot understand how to evaluate the new features contained in the external designs. Therefore, the predictor model cannot guide the generator model to fully incorporate new external features. Training the predictor model with both internal and external preferences can fully release its power, enabling it to guide the generative model. By doing so, industry practitioners can better capture consumers’ heterogeneous preferences for various styles of product designs, especially those that are not present in the existing internal designs.
@@ 6 Managerial Implications and Conclusion
Our proposed model offers a superior alternative to the prevailing industry practice that relies solely on pure GAN models without considering consumer preferences. The proposed model is cost-effective, automated, and scalable. It also generates new designs in a controllable and targeted way. The new designs are more likely to be preferred by consumers ex ante, which contrasts with the approach of generating a larger set of new designs and screening unpopular designs on an ex post basis. In this sense, our approach avoids wasting time and resources on new designs that are later shown to be less appealing.
@@
Our model provides an effective solution to the product design ideation problem. The traditional product design ideation method relies on professionally trained human designers, who have limited capacity and may not be able to generate high-quality designs on a large scale. In addition, human designers may not always understand the most recent consumer requirements and cannot systematically learn the preferences of many consumers. We show that the existing designs acquired from external sources can complement the traditional human design process. UGC images can be regarded as design ideas created by individual users in real time and serve as a rich ideation source.
@@
The proposed model offers a cost-effective alternative for obtaining consumer evaluations of new product designs. In contrast to traditional methods such as conducting consumer surveys, we leverage UGC as a source of consumer preferences. This UGC serves as a vast repository of real-time, self-evident consumer preferences that can be utilized to inform design decisions. Such information is highly relevant because it is voluntarily generated by consumers. This information is abundant, readily available online, and dynamically updated, which allows firms to keep up with constantly evolving consumer demands. UGC is especially useful for start-up firms that do not have sufficient internal resources and cannot afford the costs of designing new products using traditional methods. The use of UGC not only reduces the financial burden associated with extensive consumer surveys but also provides a more dynamic and authentic reflection of consumer tastes.
In summary, our findings suggest that industry practitioners can benefit significantly from combining preference information with generative models, moving beyond the limitations of pure GAN-based approaches. By recognizing the importance of external preferences, particularly those embedded in UGC, businesses can enhance the efficiency of their design processes and gain valuable insights into consumer preferences without conducting costly surveys.
Our method has broad applications not only for artistic designs but also for other design tasks where heterogeneous consumer preferences widely exist. Future works can study how consumer preferences can be incorporated into generative models in other contexts and evaluate the effectiveness of such approaches.
Appendix A Additional Tables and Figures
Figure A.1: Illustration from the OpenFace Project
Figure A.2: Illustration of the Reference BVLC Caffenet
Figure A.3: Histograms of the Predicted Popularity Levels (Internal Preferences)
Refer to caption Refer to caption
(a) Original Popularity Level (b) Grouped Popularity Labels
Figure A.4: Histograms of the Predicted Popularity Levels (Internal and External Preferences)
Refer to caption Refer to caption
(a) Original Popularity Level (b) Grouped Popularity Labels
Appendix B Analysis of the Traditional Chinese Theme
To further validate our results, we show that our proposed model works on a wide range of themes by applying it to generate templates belonging to the traditional Chinese theme. Similar to the graffiti case, we generate new traditional Chinese theme-based designs using the baseline DCGAN model without consumer preferences, the enhanced model where we adopt the CcGAN with internal consumer preferences, and the advanced model where we adopt the CcGAN with internal and external consumer preferences.
Similar to the graffiti theme, we gather photos from social media and image websites featuring human subjects with styles matching the internal templates of the traditional Chinese theme. These external images are then preprocessed by removing the human subjects and completing the backgrounds through inpainting, ensuring that they match the style of the internal templates offered by the company. Simultaneously, we retain the removed faces for potential later use. This screening and inpainting process yields a total of 110 traditional Chinese images. Figure B.5 provides examples of both internal and external photos that are suitable for the traditional Chinese theme. As depicted, the collected external images exhibit a similar style to that of the internal images and showcase adequate variations.
Refer to caption Refer to caption
(a) Internal images (b) External images
Figure B.5: Examples of Traditional Chinese Templates: Internal vs. External
B.1 Performance of the Prediction Model
As discussed in Section 4.1, during the training process of the prediction model on internal data, all the themes within the “main (single)” category, including the traditional Chinese theme, are utilized. Hence, there is no need to retrain Prediction Model 1, as the results obtained for both the graffiti and traditional Chinese themes are derived from the same model. The focus is on retraining Prediction Model 2, which incorporates 110 external photos as positive samples. With the addition of these external images to the 33,804 internal images, the sample size totals 33,914. Following the previous approach, 70% of the observations are allocated to the training set. To mitigate the unbalanced sample issue, the minority class is upsampled to 20,000 samples, aligning with the majority class. Prediction Model 2 is estimated using a random forest model with the same hyperparameters as those employed in the graffiti case. Table B.1 displays the accuracy of the prediction model incorporating both internal and external preferences. The results indicate that the inclusion of external preferences has a limited impact on the overall accuracy, primarily due to the restricted size of the external traditional Chinese image dataset.
Table B.1: Accuracies Achieved by the Consumer Choice Model Prediction On the Test Set
Choice Prediction Task Acc. Balanced Acc. FNR FPR
Internal Preferences Only 0.793 0.791 0.212 0.206
Internal and External Preferences 0.794 0.790 0.216 0.205
Note: This table presents the accuracy, balanced accuracy rate, false-negative rate, and false-positive rate produced for the consumer choice model prediction results on the test set. The first row presents the accuracy for the prediction model with internal preferences only, and the second row presents the accuracy for the prediction model with both internal and external preferences, where the external preferences are derived from the external traditional Chinese theme.
We then conduct a preliminary visual assessment to evaluate the effectiveness of Prediction Model 1 in terms of generating meaningful popularity labels. To do so, we identify the most- and least-preferred internal templates based on the consumer adoption rates. We then compare their styles with those of the combined internal and external templates that have the highest and lowest average predicted choice probabilities, respectively. As illustrated in Figure B.6, we observe that the external templates with high choice probabilities closely resemble the most popular internal templates. Conversely, the external templates with low choice probabilities are similar to the least-popular internal templates. This observation visually validates the accuracy of the predicted popularity labels.
Figure B.6: Comparing the Internal and External Templates by Popularity (Internal Preferences)
Most-Preferred Templates Least-Preferred Templates
Internal Templates Refer to caption Refer to caption
Combined Templates Refer to caption Refer to caption
We also conduct a preliminary visual assessment to evaluate the performance of Prediction Model 2 in terms of generating meaningful popularity labels when integrating both internal and external preferences. By comparing the most- and least-preferred internal templates with their counterparts predicted by the model (Figure B.7), we observe that the external templates with high choice probabilities exhibit styles similar to those of the most-popular internal templates, while those with low probabilities resemble the least-popular internal templates. However, in contrast to Figure 8, the similarity is less pronounced, as the new prediction model captures some external consumer preferences that are not observed in the internal data, particularly a preference for backgrounds featuring blossoms and birds. Consequently, the images with these elements are detected through the incorporation of the external preferences. Nonetheless, certain images remain consistent across both models; for example, the third images of the most-preferred external template are identical in both cases. Similarly, the third images of the least-popular external template are also consistent across models. This pattern confirms that the enhanced prediction model leverages new information derived from external preferences, enabling it to better capture consumer preferences for producing more diverse product designs.
Figure B.7: Comparing the Internal and External Templates by Popularity (Internal and External Preferences Together)
Most-Preferred Templates Least-Preferred Templates
Internal Templates Refer to caption Refer to caption
Combined Templates Refer to caption Refer to caption
B.2 Performance of the Generative Models
We first train the DCGAN model by combining the existing templates derived from internal and external sources, resulting in 119 training images for the traditional Chinese theme, 9 of which are acquired from internal data and 110 of which are acquired from external data. Since the size of our training dataset is extremely small, we again adopt the DiffAugment method (Zhao et al., (2020)), following the approach used in the graffiti case. All other hyperparameters remain consistent with those employed in the graffiti case.
Next, we proceed with the enhanced model, training the CcGAN with internal consumer preferences. Following our previous approach, we group images with similar popularity levels into bins, ensuring that the images within the same bin share an identical popularity level. Any class with fewer than 50 images is replicated to ensure a minimum of 50 images per class, thereby achieving a more balanced distribution. Subsequently, we normalize each label by dividing it by the maximum label value, resulting in 7 normalized labels ranging from 0.586 to 1. We employ the soft vicinal discriminator loss (SVDL) as the discriminator loss, with $\sigma$ and $\kappa$ selected using the rule-of-thumb formulae. The CcGAN undergoes training for 20,000 iterations with a constant learning rate of $10^{-4}$ and a batch size of 64. The latent GAN dimensionality is set at 256, and the discriminator updates twice per iteration. To address the small sample size problem, we utilize the DiffAugment method during training, as previously done in the DCGAN case. Once the CcGAN is trained, we use the generator to produce new template designs with various popularity labels, where label=0.586 represents the least-preferred class and label=1 indicates the most-preferred class.
Finally, we train the advanced model: the CcGAN incorporating both internal and external consumer preferences. Similar to the enhanced model, we utilize Prediction Model 2 to determine popularity levels for the existing templates. Continuous labels are grouped into 5 bins, with any groups containing fewer than 20 images replicated to ensure a minimum of 20 images per group. The labels are then normalized, resulting in 5 values ranging from 0.529 to 1. The SVDL is used again with the hyperparameters obtained by following the previous guidelines. After training, the generator produces new template designs with various popularity labels, ranging from least preferred (label=0.529) to most preferred (label=1).
Images generated using different models
We start by presenting the performance of the DCGAN models. Figure 8(b) compares the images generated by the trained DCGAN models from the real images belonging to the traditional Chinese theme. Figure 8(b) (a) shows the real traditional Chinese images, where these templates feature a light-colored background with some plum blossoms. As shown in Figure 8(b) (b), the styles of the DCGAN-generated images are similar to those of the real images. Significant variations are also observed in the styles of these generated images. Therefore, by employing a simple visual examination, we know that the DCGAN model does a good job when generating traditional Chinese theme templates.
Figure B.8: Real vs. DCGAN-Generated Images Belonging to the Traditional Chinese Theme
(a) Real Images
(b) DCGAN-Generated Images
Next, we show that the CcGAN models can controllably generate new designs with different popularity levels. Figure B.9 shows the generated images, ranging from the least-preferred class to the most-preferred class. The top row displays the images generated using internal preferences in the CcGAN models, while the bottom row shows the images generated with both internal and external preferences. A clear trend emerges: as the label increases, the style of the generated images gradually shifts, indicating alignment with the consumer preferences across different classes.
Figure B.9: Traditional Chinese Images Generated with Continuous Labels
B.3 Comparing Different GAN Models
We now compare the results of these three methods and check whether any positive effect results from incorporating consumer preferences into the process of generating new designs. Table B.2 presents the choice probabilities predicted for the generated images using various methods. The first column is normalized to 1, representing the choice probability of the real images, while the subsequent columns display their percentage improvements relative to the first column. The first row illustrates the choice probabilities predicted when only internal preferences are incorporated into the prediction model, while the second row depicts the choice probabilities predicted when both internal and external preferences are considered. The table compares the DCGAN, the CcGAN with internal preferences, and the CcGAN with both internal and external preferences. The DCGAN relies solely on external templates, resulting in a decrease in performance (-4.55%) compared to that achieved with internal real images. This decline is attributed to the inability of the predictor model to recognize potentially appealing features from external templates. The CcGAN with internal preferences yields improved performance by guiding a GAN with internal preferences, yet its efficacy is constrained because the predictor model fails to train on external features. Incorporating external preference information enhances the performance of a GAN by leveraging the appealing features acquired from external templates. The predictor model learns the preferred features from external consumers, assuming that the preferences of internal and external consumers are similar. Consequently, the predictor-guided CcGAN provides the largest improvement (11.35%), generating templates with popular features from both sources. The performance of the DCGAN also improves (8.97%) as the predictor model recognizes popular external features in the generated templates.
Table B.2: Predicted Choice Probabilities for the Generated Images
Methods Internal Real Images DCGAN CcGAN (Label=1)
Internal Preferences Only 1.00 -4.55% -0.84%
Internal + External Preferences 1.00 8.97% 11.35%
Note: This table presents the average selection probabilities predicted for the images generated using different methods. We normalize the first column to 1 (i.e., make the selection probability for real images=1) and present the other columns as percentage improvements relative to the first column. The first row calculates the selection probabilities for the internal real images, DCGAN-generated images, and CcGAN-generated images (with internal preferences incorporated) for the most-preferred class (label=1) using the prediction model where only internal preferences are added. The second row calculates the selection probabilities for the internal real images, DCGAN-generated images, and CcGAN-generated images (with internal and external preferences incorporated) for the most-preferred class (label=1) using the prediction model with both internal and external preferences added.

### v3

@@ Introduction
Notes: This table presents the average choice probabilities predicted for real images and images generated using different GAN models. The first (second) row shows the choice probabilities predicted by Prediction Model 1 (Prediction Model 2). Choice probabilities of real internal images are normalized to 1 in the first column, with percentage changes shown in other columns.
Personalized Design Generation
Our proposed framework is well suited for personalized design generation. The CcGAN with both internal and external preferences is trained based on population-wide popularity labels (i.e., the average choice probabilities across consumers). We can use this general CcGAN as a pre-trained model and re-train it using personalized popularity labels to obtain a personalized CcGAN. For a target consumer, we first use the trained predictor to calculate his choice probabilities for each input image, call these probabilities the personalized popularity labels. We then use these personalized labels to guide the re-training of the generative model. The re-trained generative model is now able to conduct personalized design generation by directionally generating new images that is preferred by this specific consumer.
We randomly choose one of the existing internal consumers and conduct a personalization exercise for him. Figure 11(a) shows the target consumer and the photo he took with the chosen internal template. We use the trained predictor model to calculate his choice probabilities for each input internal and external templates. We then use these probabilities as personalized popularity labels to re-train the general CcGAN for 8000 iterations. This gives us the personalized CcGAN for this consumer. We instruct the personalized CcGAN to generate new templates from the most preferred class.
Figure 11: Personalized Design Generation
To evaluate the performance of the generated personalized templates, we define a personalized distance-based metric by calculating the distance of each generated template to a set of target images that represent the target consumer’s most- and least-preferred existing templates. The calculation procedure is similar to the previous distance-based metric test, but the target images are redefined to be personalized: 1) the internal template chosen by the target consumer, representing his most-preferred internal template; 2) the internal template chosen by the internal consumer who looks the least similar to him (based on the distance of their facial appearance embeddings), representing his least-preferred internal template; 3) the external template chosen by the most similar external consumer, representing his most-preferred external template; and 4) the external template chosen by the least similar external consumer, representing his least-preferred external template. Figure 11(b) displays the four target images. Intuitively, consumers who resemble the focal consumer are likely to share similar preferences, making their choices more appealing to him. Indeed, we find that his most-preferred templates share similar patterns, which are different from his least-preferred templates. The personalized distance-based metric measures how similar the personalized templates are to his most- and least-preferred templates. A smaller metric represents better performance.
We use the trained personalized CcGAN to generate 100 personalized designs. Figure 11(c) shows sample personalized designs. Table 5 presents the means and medians of the personalized distance-based metrics by the personalized and non-personalized models. We find that the personalized CcGAN with both preferences performs the best, followed by the general CcGAN with both preferences, and the DCGAN performs the worst. The results suggest that incorporating personalized preference yields even better designs than incorporating population-level preference in terms of generating designs that are attractive to a specific consumer.
Table 5: Personalized Distance-Based Metrics: Mean and Median
@@
Statistics DCGAN CcGAN with internal and external preferences CcGAN with internal and external preferences (personalized)
Mean 1.22 1.10 0.97
Median 1.22 1.05 0.96
@@
Note: This table presents the mean and median distance-based metrics for images generated by different GAN models. Smaller values indicate the generated templates are visually similar to popular internal templates and distinct from unpopular ones, suggesting higher consumer popularity.
Conclusion
In this paper, we propose a novel framework to transform the product design process in the era of generative AI. The unique contribution of our work is that our proposed model systematically incorporates consumer preferences and external UGC into the design generation process. Two important managerial insights emerge from our results. First, it is beneficial to incorporate consumer preference information into the generation process because it can guide the generator to directionally generate consumer-preferred designs. Second, it is not sufficient to incorporate only external designs; external preference information is also required for the generative model to fully incorporate new appealing features from external designs.
@@ Conclusion
Our framework offers a superior alternative to the prevailing industry practice that relies solely on pure generative models without considering consumer preferences. The proposed model is preference-aware, cost-effective, automated, and scalable, generating new designs in a controllable and targeted way. The new designs are more likely to be preferred by consumers ex ante, unlike the traditional method of creating many designs and discarding unpopular ones later. Our approach saves resources by avoiding less appealing designs.
@@
Our framework provides an effective solution to the product design ideation problem. Traditional methods depend heavily on professionally trained human designers, who have limited capacity and may struggle to produce high-quality designs on a large scale. Additionally, these designers might not always be fully attuned to the latest consumer demands and often cannot systematically learn the preferences of a broad audience. Our approach enhances the traditional design process by incorporating existing designs from external sources, which can complement human efforts. User-generated content (UGC) images, in particular, can be viewed as real-time design ideas created by individual users, offering a rich and diverse source for ideation.
@@
Our framework provides a cost-effective way to evaluate new product designs by using user-generated content (UGC) instead of traditional consumer surveys. UGC offers a vast, real-time repository of consumer preferences that inform design decisions. This data is highly relevant as it is voluntarily generated by consumers, abundant, readily available online, and constantly updated, helping firms stay aligned with evolving consumer demands. UGC is particularly beneficial for start-ups lacking resources for costly traditional design methods, reducing financial burdens and offering a dynamic, authentic reflection of consumer tastes.
In conclusion, our findings highlight the benefits for industry practitioners of integrating consumer preferences and external UGC into generative models. Unlike pure GAN-based methods, our approach utilizes external preference data from UGC to enhance design processes and extract valuable consumer insights without costly surveys. This framework is widely applicable in artistic, creative, and other domains with diverse consumer preferences. Future research could investigate embedding consumer preferences into generative models across various contexts, expanding understanding of the effectiveness and scalability of these approaches.